OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1).

OpenStudy (amistre64):

2, 1 --3, -2 ------- 5,-1 slope = -1/5 y-1 = -1/5 (x-2)

OpenStudy (aroub):

y=mx+b y=1 1=2(-1/5)+b 1=-2/5+b solve the equation b=find it out :)

OpenStudy (amistre64):

point slope form; but yes, that is a good way to determing the slope intercept form

OpenStudy (aroub):

yes their is 2 ways to do it

OpenStudy (amistre64):

no; there is only one way to use the point slope form of a line. There are many ways describe the equation of a line. .... but only one way to describe it in point slope form....

OpenStudy (aroub):

yes but their is y=mx+b and y-\[_{?}\]=m(x-\[_{?}\]

OpenStudy (aroub):

the y intercept form and the point slope formula u can do it both ways

OpenStudy (amistre64):

"Choose the point-slope form of the equation": y - y1 = m(x-x1) ; where point is P(x1,y1). Now if you want to do this the long way..... then yes, try to fight thru the slope-intercept form to get to it. By the way, I have never really found a good use for the point-slope form, unless we are looking for the equation of a tangent line to the curve, then its pretty handy :)

OpenStudy (aroub):

hahaaha yes it is :) which grade (or class or year )are u ???

OpenStudy (amistre64):

im 39 years old; i dont think I have a grade any more ;)

OpenStudy (aroub):

hahahaha ya i dont think so lol sry for asking do u teach or something???

OpenStudy (amistre64):

i just help people online with their math problems if i can :) im going for my masters degree in math to be a college math teacher at the moment tho.

OpenStudy (aroub):

oh! ya really very nice:)

OpenStudy (aroub):

in which country??

OpenStudy (amistre64):

here in the usa...